I need help with another problem please. my solution doesn't seemt right.
solving for b^2, a^2+b^2=c^2. A=9^2 (81), b=unknown, C=square root of 17. I come up with b^2=-64. Is that correct? Thanks

Answer by stanbon(75887)

solving for b^2, a^2+b^2=c^2. A=9^2 (81), b=unknown, C=square root of 17
b^2=c^2-a^2
b^2=(sqrt17)^2-(81)
b^2=17-81
b^2=-64
Cheers,
Stan H.